OpenStudy (peterg2001):

help please will medal Find the range of the given data set. -4, -3, -1, -1, 0, 1 -5 -3 3 5

OpenStudy (peterg2001):

@haleyelizabeth2017

OpenStudy (haleyelizabeth2017):

My bad, I was thinking of something else. What is 1-(-4)? hint: -(-4)=+4

OpenStudy (anonymous):

Range = Largest item - Smallest item = 1 - (- 4) = 1 + 4 = 5 Hope this helps :)